# Can LLM-Generated Textual Explanations Enhance Model Classification Performance? An Empirical Study

### v1

@@ Abstract
Keywords:
Explainable natural language processing Natural language explanations Natural language inference Large language models.
@@ 1 Introduction
Recent NLP advancements are driven by PLMs and LLMs, achieving state-of-the-art results across various tasks [2]. However, their black-box nature limits understanding of their predictions, prompting increased interest in Explainable NLP, where methods from Explainable AI explain model decision-making [31] to enhance trust and transparency, which is essential for advancing practical applications in sensitive domains.
A key challenge in Explainable NLP is the lack of definitive ground-truth explanations [17]. Researchers address this by collecting human-generated textual explanations, creating explainable datasets [6]. These datasets serve both as benchmarks for evaluating model-generated explanations and as training data to improve models’ predictive performance [38]. However, human annotation is resource-intensive, impacting dataset scale and quality [29]. Recently, leveraging LLMs’ text-generation capabilities for explanations has gained attention [37], though evaluating these explanations’ quality and effectiveness in downstream tasks remains an open research question.
@@ 2 Background and Related Work
Natural Language Inference (NLI) is one of the most fundamental NLP tasks [9]. The goal is, given two pieces of text, a premise and a hypothesis, to determine a logical relation between them as one of the three classes: entailment, contradiction, or neutral. The turning point in NLI was the construction of the Stanford NLI (SNLI) corpus in 2015 [1], a dataset of half a million examples, constructed with crowd-sourced effort where photos were captioned and then paired with entailed, contradicted, or neutral sentences written by annotators. Modern PLMs like BERT [5] and RoBERTa [20], as well as autoregressive LLMs like GPT, can often solve popular NLI datasets with an above-human performance, owing to the linguistic patterns and world knowledge acquired during their pre-training on huge corpora.
@@
Explainable NLP and Datasets The growing interest in Explainable NLP is evident from multiple surveys like [21, 38], some addressing specific tasks or methods [22]. Interest in explainable NLP has led to the creation of explainable datasets for tasks such as hate-speech classification [23] and claim verification [33]. A comprehensive review of these datasets is provided in [38]. Textual explanations typically fall into highlights, structured, or free-text (natural language) categories and are annotated by authors, experts, and crowd-sourcing with most datasets relying on human annotators. However, human annotation presents several challenges. Collecting high-quality explanations is time-consuming and resource-intensive [10].Human annotators’ explanations may also suffer from subjectivity and inconsistency, potentially hindering model performance rather than aiding it [39]. Additionally, the diversity in explanation types introduces further complexities [32].
@@
Generating LLM-explanations Due to the limitations of human-annotated explanations, recent research has explored using LLMs to generate Natural Language Explanations (NLE) and justifications for model decisions. Compared to traditional post-hoc feature attribution methods, NLEs provide human-readable justifications, which can enhance transparency and user understanding. [24] employed LLMs as rationalizers for knowledge-intensive tasks such as multiple-choice question answering. [35] investigated improving LLM-generated NLE quality through a tandem learning setup. [40] examined how various prompting techniques, such as CoT, can improve NLEs on commonsense reasoning tasks. These studies demonstrate the growing interest in leveraging LLMs to generate and refine NLEs, particularly for tasks requiring explanation-driven reasoning. However, none of the previously mentioned works investigate how extending datasets with LLM-generated explanations can impact the performance of PLMs and LLMs on downstream tasks.
@@
Evaluating NLEs NLEs are text snippets and can be evaluated with standard NLG metrics [30]. When human-written (gold) references exist, reference-based metrics are applicable. Traditional metrics, such as BLEU [25] and ROUGE [18], assess word overlaps between generated and reference texts. However, these metrics have become less suitable with the rise of LLMs, as they penalize expressive variations in wording. Consequently, semantic metrics like the embedding-based BERTScore [41] and distribution-based MAUVE [27] have gained popularity. Recently, evaluation methods using LLM-as-judge metrics have emerged, employing crafted prompts for LLMs to return numerical scores assessing generated texts, exemplified by G-Eval [19].
@@
Closely related work Among the previously mentioned works, the closest to ours is [39], which investigates how human explanations can impact the predictions of two PLMs. However, their study is limited to BART and T5 and focuses solely on human explanations. Additionally, [11] reviews studies employing different types of human explanations (highlights, structured, and free-text) to improve NLP models. However, they solely review studies incorporating human-annotated explanations. In contrast, while we also incorporate human explanations, our primary focus is on generating and investigating LLM-generated NLEs. We evaluate the impact of these explanations on four PLMs, including the recent ModernBERT [36], as well as three LLMs of varying sizes.
@@ 3.1 Datasets
We use two datasets in our experiments. The first dataset, e-SNLI [3], is an extension of the SNLI dataset with human-annotated natural language explanations. It contains premise-hypothesis pairs labeled as entailment, neutral, or contradiction, depending on how the premise relates to the hypothesis. The second dataset is the HealthFC dataset [34]. It consists of 750 health-related claims, labeled by medical experts and backed with evidence from systematic reviews and clinical trials. Each claim is paired with pieces of evidence and includes a verdict (supported, refuted, not enough information), as well as brief explanations for the verdict. For our experiments, we extracted a balanced subset of e-SNLI consisting of 840 examples, ensuring an equal representation of entailment, neutral, and contradiction instances. This subset size was deliberately chosen to closely match the 750-instance HealthFC dataset, enabling a fair and controlled comparison across our evaluation framework. Even though HealthFC is officially a dataset for automated fact-checking (claim verification), it is common to model this task as an NLI task. We provide further details on the datasets with examples in Appendix 5.1.
@@ 3.2 Generating Natural Language Explanations with LLMs
In our pipeline, we focus on generating NLEs using multiple LLMs. We further extend both datasets we consider with explanations we generate using GPT-4o mini [13], Mixtral-7B [14], Gemma2-9B [7], and LLama3-70B [8]. For Mixtral-7B, Gemma2-9B, and LLama3-70B we use the APIs provided by Groq [^2] https://groq.com/ while for GPT-4o mini we use OpenAI APIs [^3] https://platform.openai.com/. We selected LLMs ranging in size from 7B to 70B parameters444along with GPT-4o mini, whose exact size is unknown., to analyze how these factors influence both the quality of the generated text and the impact of generated explanations on downstream task performance. The rationale for selecting diverse LLMs, rather than models within the same family differing only in size, is to ensure a broader variety in the sources of explanations. We discuss later in the paper how this approach could be expanded in future work.
@@ 3.3 Evaluating LLM-Natural Language Explanations
We compare LLM-generated explanations with human-provided explanations from our selected datasets . Specifically, we employ the widely used BLEU, ROUGE, and BERTScore metrics. Beyond these conventional metrics, we incorporate the recent MAUVE. and the LLM-as-judge G-Eval framework that has been increasingly used in recent NLG research. We use G-Eval to measure human likeness in LLM-generated explanations, in particular, the clarity, coherence, and structure of the LLM-generated explanation. We refer readers to Appendix 5.4 for further details on G-Eval computation, implementation specifics, and metric libraries.
@@ 3.4 Models for NLI Predictions
Fine-tuning PLMs. For the downstream NLI task predictions, we use four PLMs (BERT , DeBERTa [12] , RoBERTa , and ModernBERT [36]). For each run with a certain kind of or without explanations, we perform a 80/20 train/test split and fine-tune the PLMs on the train set for 10 epochs using the AdamW optimizer with a learning rate of 3e-6 for ModernBERT and 1e-5 for the other PLMs. We repeat this five times with a stratified 5-fold cross-validation and report results averaged over the five splits.
@@ 5 Conclusion
Future work will explore extending the framework to a broader set of datasets to encompass a wider range of tasks and complexities and further refine prompt engineering and explanation generation via refinement techniques [35] , verification and refinement [28], and consistency fine-tuning [4]. Additionally, incorporating emerging evaluation metrics such as TIGERScore [15] and Prometheus [16] will enable more comprehensive quality assessments, while comparisons with advanced reasoning LLMs like OpenAI o3 and DeepSeek R1 could further validate our approach. In addition, we plan to extend our selection of LLMs used for generating explanations by experimenting with LLMs from the same family of different sizes (e.g., Gemma-9b vs Gemma-27b) to measure the impact of size on the quality of explanations per the metrics used in this study. Finally, another point of improvement is measuring and improving the faithfulness of self-explanations by LLMs [26], as we have observed that when asked to output the most important words for their predictions, LLMs frequently assign high importance to peripheral words in the prompt such as those describing the labels or denoting parts of the input such as the explanations provided.

### v2

@@ 1 Introduction
Recent NLP advancements are driven by PLMs and LLMs, achieving state-of-the-art results across various tasks [brown2020language]. However, their black-box nature limits understanding of their predictions, prompting increased interest in Explainable NLP, where methods from Explainable AI explain model decision-making [soegaard-2022-xnlp-book] to enhance trust and transparency, which is essential for advancing practical applications in sensitive domains.
A key challenge in Explainable NLP is the lack of definitive ground-truth explanations [lei2016rationalizing]. Researchers address this by collecting human-generated textual explanations, creating explainable datasets [deyoung2020eraser]. These datasets serve both as benchmarks for evaluating model-generated explanations and as training data to improve models’ predictive performance [wiegreffe2020annotated]. However, human annotation is resource-intensive, impacting dataset scale and quality [rajani2019explain]. Recently, leveraging LLMs’ text-generation capabilities for explanations has gained attention [wei2022chain], though evaluating these explanations’ quality and effectiveness in downstream tasks remains an open research question.
@@ 2 Background and Related Work
Natural Language Inference (NLI) is one of the most fundamental NLP tasks [gubelmann2024capturing]. The goal is, given two pieces of text, a premise and a hypothesis, to determine a logical relation between them as one of the three classes: entailment, contradiction, or neutral. The turning point in NLI was the construction of the Stanford NLI (SNLI) corpus in 2015 [bowman-etal-2015-large], a dataset of half a million examples, constructed with crowd-sourced effort where photos were captioned and then paired with entailed, contradicted, or neutral sentences written by annotators. Modern PLMs like BERT [devlin2019bert] and RoBERTa [liu2019roberta], as well as autoregressive LLMs like GPT, can often solve popular NLI datasets with an above-human performance, owing to the linguistic patterns and world knowledge acquired during their pre-training on huge corpora.
@@
Explainable NLP and Datasets The growing interest in Explainable NLP is evident from multiple surveys like [madsen-xnlp-survey-2023, wiegreffe2020annotated], some addressing specific tasks or methods [Mardaoui-2021-lime-survey]. Interest in explainable NLP has led to the creation of explainable datasets for tasks such as hate-speech classification [mathew2021hatexplain] and claim verification [vladika-etal-2025-step]. A comprehensive review of these datasets is provided in [wiegreffe2020annotated]. Textual explanations typically fall into highlights, structured, or free-text (natural language) categories and are annotated by authors, experts, and crowd-sourcing with most datasets relying on human annotators. However, human annotation presents several challenges. Collecting high-quality explanations is time-consuming and resource-intensive [hartmann-2022-survey-human-explanation-performance]. Human annotators’ explanations may also suffer from subjectivity and inconsistency, potentially hindering model performance rather than aiding it [yao-2023-human-explanations-helpful]. Additionally, the diversity in explanation types introduces further complexities [tan2021diversity].
@@
Generating LLM-explanations Due to the limitations of human-annotated explanations, recent research has explored using LLMs to generate Natural Language Explanations (NLE) and justifications for model decisions. Compared to traditional post-hoc feature attribution methods, NLEs provide human-readable justifications, which can enhance transparency and user understanding. [mishra-etal-2024-characterizing-rationalizers] employed LLMs as rationalizers for knowledge-intensive tasks such as multiple-choice question answering. [wang-etal-2025-cross-refine] investigated improving LLM-generated NLE quality through a tandem learning setup. [wei-jie-etal-2024-interpretable-reasoning-nle] examined how various prompting techniques, such as CoT, can improve NLEs on commonsense reasoning tasks. These studies demonstrate the growing interest in leveraging LLMs to generate and refine NLEs, particularly for tasks requiring explanation-driven reasoning. However, none of the previously mentioned works investigate how extending datasets with LLM-generated explanations can impact the performance of PLMs and LLMs on downstream tasks.
@@
Evaluating NLEs NLEs are text snippets and can be evaluated with standard NLG metrics [schmidtova-etal-2024-automatic-metrics]. When human-written (gold) references exist, reference-based metrics are applicable. Traditional metrics, such as BLEU [papineni-etal-2002-bleu] and ROUGE [lin-2004-rouge], assess word overlaps between generated and reference texts. However, these metrics have become less suitable with the rise of LLMs, as they penalize expressive variations in wording. Consequently, semantic metrics like the embedding-based BERTScore [Zhang2020BERTScore] and distribution-based MAUVE [pillutla2021mauve] have gained popularity. Recently, evaluation methods using LLM-as-judge metrics have emerged, employing crafted prompts for LLMs to return numerical scores assessing generated texts, exemplified by G-Eval [liu-etal-2023-G-eval].
@@
Closely related work Among the previously mentioned works, the closest to ours is [yao-2023-human-explanations-helpful], which investigates how human explanations can impact the predictions of two PLMs. However, their study is limited to BART and T5 and focuses solely on human explanations. Additionally, [hartmann-2022-survey-human-explanation-performance] reviews studies employing different types of human explanations (highlights, structured, and free-text) to improve NLP models. However, they solely review studies incorporating human-annotated explanations. In contrast, while we also incorporate human explanations, our primary focus is on generating and investigating LLM-generated NLEs. We evaluate the impact of these explanations on four PLMs, including the recent ModernBERT [warner2024smarter], as well as three LLMs of varying sizes.
@@ 3.1 Datasets
We use two datasets in our experiments. The first dataset, e-SNLI [camburu2018e], is an extension of the SNLI dataset with human-annotated natural language explanations. It contains premise-hypothesis pairs labeled as entailment, neutral, or contradiction, depending on how the premise relates to the hypothesis. The second dataset is the HealthFC dataset [vladika-etal-2024-healthfc]. It consists of 750 health-related claims, labeled by medical experts and backed with evidence from systematic reviews and clinical trials. Each claim is paired with pieces of evidence and includes a verdict (supported, refuted, not enough information), as well as brief explanations for the verdict. For our experiments, we extracted a balanced subset of e-SNLI consisting of 840 examples, ensuring an equal representation of entailment, neutral, and contradiction instances. This subset size was deliberately chosen to closely match the 750-instance HealthFC dataset, enabling a fair and controlled comparison across our evaluation framework. Even though HealthFC is officially a dataset for automated fact-checking (claim verification), it is common to model this task as an NLI task. We provide further details on the datasets with examples in Appendix 5.1.
@@ 3.2 Generating Natural Language Explanations with LLMs
In our pipeline, we focus on generating NLEs using multiple LLMs. We further extend both datasets we consider with explanations we generate using GPT-4o mini [hurst2024gpt], Mixtral-7B [jiang2024mixtral], Gemma2-9B [team2024gemma], and LLama3-70B [dubey2024llama]. For Mixtral-7B, Gemma2-9B, and LLama3-70B we use the APIs provided by Groq [^2] https://groq.com/ while for GPT-4o mini we use OpenAI APIs [^3] https://platform.openai.com/. We selected LLMs ranging in size from 7B to 70B parameters444along with GPT-4o mini, whose exact size is unknown., to analyze how these factors influence both the quality of the generated text and the impact of generated explanations on downstream task performance. The rationale for selecting diverse LLMs, rather than models within the same family differing only in size, is to ensure a broader variety in the sources of explanations. We discuss later in the paper how this approach could be expanded in future work.
@@ 3.3 Evaluating LLM-Natural Language Explanations
As we focus on generating natural language explanations, we evaluate their quality using some of the widely adopted metrics in NLG research we described in Section 2. We compare LLM-generated explanations with human-provided explanations from our selected datasets. Specifically, we employ the widely used BLEU, ROUGE, and BERTScore metrics. Beyond these conventional metrics, we incorporate the recent MAUVE, a distribution-based metric that quantifies the divergence between generated and human-written texts using Kullback–Leibler (KL) divergences in a quantized embedding space and also the LLM-as-judge G-Eval framework that has been increasingly used in recent NLG research. We use G-Eval to measure human likeness in LLM-generated explanations, in particular, the clarity, coherence, and structure of the LLM-generated explanation. We refer readers to Appendix 5.4 for further details on G-Eval computation, implementation specifics, and metric libraries.
@@ 3.4 Models for NLI Predictions
Fine-tuning PLMs. For the downstream NLI task predictions, we use four PLMs (BERT , DeBERTa [he2020deberta] , RoBERTa , and ModernBERT [warner2024smarter]). For each run with a certain kind of or without explanations, we perform a 80/20 train/test split and fine-tune the PLMs on the train set for 10 epochs using the AdamW optimizer with a learning rate of 3e-6 for ModernBERT and 1e-5 for the other PLMs. We repeat this five times with a stratified 5-fold cross-validation and report results averaged over the five splits.
@@ 5 Conclusion
Future work will explore extending the framework to a broader set of datasets to encompass a wider range of tasks and complexities and further refine prompt engineering and explanation generation via refinement techniques [wang-etal-2025-cross-refine] , verification and refinement [quan-etal-2024-NLEs-refinement], and consistency fine-tuning [chen-etal-2025-EC-Fine-tune]. Additionally, incorporating emerging evaluation metrics such as TIGERScore [jiang2024tigerscore] and Prometheus [kim2024prometheus] will enable more comprehensive quality assessments, while comparisons with advanced reasoning LLMs like OpenAI o3 and DeepSeek R1 could further validate our approach. In addition, we plan to extend our selection of LLMs used for generating explanations by experimenting with LLMs from the same family of different sizes (e.g., Gemma-9b vs Gemma-27b) to measure the impact of size on the quality of explanations per the metrics used in this study. Finally, another point of improvement is measuring and improving the faithfulness of self-explanations by LLMs [parcalabescu2023measuring], as we have observed that when asked to output the most important words for their predictions, LLMs frequently assign high importance to peripheral words in the prompt such as those describing the labels or denoting parts of the input such as the explanations provided.